A man is allowed 25 shots at a target. For every bull's-eye he receives $5.00, but for every miss he pays $2.50. How many bull's-eyes did he achieve if he made a profit of $5.00?
Answer by checkley79(3341)

5x-2.50(25-x)=5
5x-62.5+2.5x=5
7.5x=5+62.5
7.5x=67.5
x=67.5/7.5
x=9 the number of bull's-eye.
25-9=16 misses.
Proof:
5*9-2.5*16=5
45-40=5
5=5